# It’s Not Easy Being Green: On the Energy Efficiency of Programming Languages

### v1

@@ 4.2.2. Garbage Collection
An important property of higher-level languages and their implementations is garbage collection, which imposes significant performance overhead (Hertz and Berger, 2005). Out of the 13 language implementations discussed in this paper, 10 use a garbage collector at runtime. Garbage collectors can typically be configured and tuned for each application to offer maximal performance, which is standard practice in real-world applications to optimize for application throughput and/or latency. For example, Go’s garbage collector can be disabled entirely with the GOGC environment variable. When doing so, we observe a 2.8×2.8$\times$2.8 × speedup on the allocation-intensive binary-trees benchmark.
@@ 4.3. Decomposing Energy into Power and Time
Figure 8. Energy consumption is the product of power and time (§4.3). Splitting into these two components allows separate consideration of each factor. As Figure 9 shows, there is no significant difference in power draw between implementations when we keep external factors constant, hence the impact of programming language implementation on energy consumption beyond runtime is negligible.
@@
Figure 9. Average power draw (package (PKG) + DRAM) does not differ significantly across programming language implementations (§4.3). Each point represents a single benchmark implementation (i.e., a single language implementation / benchmark pair). Points exhibit extremely low variance, and so are overlapped in the graph. After controlling for external factors such as varying number of active cores and runtime, average power draw is constant across benchmark and programming languages implementations. The red line indicates the mean power draw of $189.8\pm 0.5$ W.
@@
Fundamentally, energy consumption is the product of power and time. Figure 8 decomposes these two factors. To test if the choice of programming language implementation has an effect on power draw beyond runtime, we fix other factors that may affect power draw. We divide by execution time to normalize for runtime differences. We limit the program to a single core via Docker’s --cpuset-cpus argument. To ensure frequency scaling and throttling do not increase power draw variance in our measurements, we pin CPUs to their minimum frequency with turbo mode disabled. Figure 9 shows that there is no significant difference in power draw between programming language implementations once these outside factors are controlled for.
In theory, any language implementation with a runtime system running in parallel may affect energy consumption beyond the increased runtime, since increased core usage will increase power draw, as Section 4.4.1 details below. This hidden increased parallelism could occur with a parallel garbage collector or just-in-time compilation threads executing alongside the main program. In practice, compilation threads are primarily active only during startup, and both are rarely significant compared to the main program’s execution and other general overhead costs. For this reason, we draw this edge in gray in our model.
@@ 4.4. Contributors to Power Draw
(a) Average package (PKG) power draw by average number of active cores (§4.4.1). Log regression: $y=30\log_{2}x+242\left(R^{2}=0.70\right)$.
@@
(b) Average DRAM power draw by memory activity (§4.4.2), approximated using LLC misses. Linear regression: $y=4.47\cdot 10^{-8}x+11.6\left(R^{2}=0.94\right)$.
@@
Figure 11. Power draw is much more significantly affected by the number of active cores than memory activity (§4.4). Each point on those graphs represents a single benchmark implementation (i.e., a single language implementation / benchmark pair).
@@ 4.4.1. Number of Active Cores
Figure 11(a) shows the relationship between the number of active cores and power draw. Using multiple cores, energy usage grows up to around 2×2$\times$2 × the base power draw compared to using a single core. We fit a log curve to the data $\left(R^{2}=0.70\right)$, as the relationship does not appear linear, but results may vary across platforms. Denoting $\mbox{Power}\left(x\right)$ as the average power draw using $x$ cores, we find that, on our system, doubling the number of cores used increases average power draw by only roughly 30W:
$\mbox{Power}\left(x\right)$ $=30\log_{2}x+242$
$\Rightarrow\mbox{Power}\left(2x\right)$ $=30\log_{2}\left(2x\right)+242$
$\Rightarrow\mbox{Power}\left(2x\right)$ $=30\log_{2}2+30\log_{2}x+242$
$\Rightarrow\mbox{Power}\left(2x\right)$ $=\mbox{Power}\left(x\right)+30$
@@
$\frac{\mbox{Power}\left(2x\right)}{\mbox{Power}\left(x\right)}=\frac{30\log_{2}\left(2x\right)+242}{30\log_{2}x+242}\leq 113\%\quad\left(1\leq x\leq 128\right)$
@@ 4.4.2. Memory Activity
Previous research has shown that roughly 40% of dynamic random access memory (DRAM) energy consumption is constant and required to refresh memory cells to keep the system running properly, while the remaining 60% is related to read and write activity (Seol et al., 2016). Section 3.2.2 introduces last level cache (LLC) misses as a better proxy for memory activity: any read or write request that cannot be fulfilled by the LLC will have to go to DRAM. Figure 11(b) shows a strong linear correlation between LLC misses and increased power draw $\left(R^{2}=0.94\right)$.
@@ 4.5.1. Concurrency
Further, since TypeScript is a strict superset of JavaScript and we test them using the same runtime, there should be no differences in runtime or energy consumption. With some minor edits in four benchmarks, all JavaScript benchmark implementations pass the TypeScript compiler without any errors, and yield equal performance.
@@ 4.5.2. Third-Party Libraries
Using third-party libraries also allows for easily crossing language boundaries, for instance to access lower-level system utilities or to obtain greater performance. In Python, using NumPy can yield up to 60,000×60,000$\times$60 , 000 × faster code when multiplying matrices (Leiserson et al., 2020; Berger et al., 2023). To some degree, this has even become the “correct” way of writing Python code, by using low-level languages for computation intensive tasks and wrapping those in Python functions. Even C++ may cross over to code originally written using C, Fortran, or direct Assembly for performance when using BLAS.
@@ 5. Related Work
Previous sections, notably Section 2, discuss the main line of research relating programming languages and energy consumption (Couto et al., 2017; Pereira et al., 2017, 2021). These studies discuss and rank programming languages based on runtime, energy consumption, and memory usage. Contrary to their results, we show in this paper that the interplay between these properties is complex and depends on many factors not taken into account in their work.
The processor’s operating voltage and frequency also have an impact on energy consumption. This impact has been studied in the context of trading off speed for reduced energy consumption, where compiling for speed yields better energy efficiency in the general case (Yuki and Rajopadhye, 2013). For some specific applications, such as sparse matrix computations, compile-time techniques may decrease energy usage without impact to execution time by leveraging load imbalance (Chen et al., 2005). Other work has shown that using characteristics of a program at runtime to reduce the processor’s voltage or frequency can yield higher energy efficiency for a comparably smaller performance degradation (Hsu and Feng, 2005). For example, performance counter information can be used to scale the processor’s frequency (Weissel and Bellosa, 2002). These approaches are orthogonal to the choice of programming language and could be applied in many programming environments.
Previous research also addresses the energy consumption of different data structures in certain languages, notably Java collections (Hasan et al., 2016; Pereira et al., 2016; Oliveira et al., 2021). Data structures can be interesting candidates to study with an energy-focused lens as they can have a substantial impact on locality and memory activity, which are factors of energy consumption (as Section 4.4.2 describes). In many cases, however, programmers only have a limited choice of data structures that would be a good fit for their use case or algorithm.
Finally, past work has also investigated energy efficiency of concurrency related structures, such as Haskell’s data sharing primitives (Lima et al., 2016) or Java’s thread management constructs (Pinto et al., 2014). As Section 4.4.1 argues, these studies confirm that the power usage patterns of the processor and the machine as a whole are more complex when parallel execution is involved.
@@ 6. Conclusion
Using this causal model, we investigate and explain anomalies in previous research, finding that many factors such as parallelism level, benchmark implementation specifics, or language implementation properties must be taken into account for a fair comparison. Our results suggest that the choice of programming language has no significant impact on energy consumption beyond runtime. Programmers aiming to reduce energy consumption can do so even in “inefficient” programming languages like Python by using faster language implementations, employing faster algorithms, and using fast third-party (native) libraries.

### v2

@@ 4.2.2. Garbage Collection
An important property of higher-level languages and their implementations is garbage collection, which imposes significant performance overhead (Hertz and Berger, 2005). Out of the 13 language implementations discussed in this paper, 10 use a garbage collector. Garbage collectors can typically be configured and tuned for each application to offer maximal performance, which is standard practice in real-world applications to optimize for application throughput and/or latency. For example, Go’s garbage collector can be disabled entirely with the GOGC environment variable. When doing so, we observe a 2.8×2.8$\times$2.8 × speedup on the allocation-intensive binary-trees benchmark.
@@ 4.3. Decomposing Energy into Power and Time
Figure 9. Energy consumption is the product of power and time (§4.3). Splitting into these two components allows separate consideration of each factor. As Figure 10 shows, there is no significant difference in power draw between implementations when we keep external factors constant, hence the impact of programming language implementation on energy consumption beyond execution time is negligible.
@@
Figure 10. Average power draw (package (PKG) + DRAM) does not differ significantly across programming language implementations (§4.3). Each point represents a single benchmark run. Points exhibit extremely low variance, and so are overlapped in the graph. After controlling for external factors such as varying number of active cores and execution time, average power draw is constant across benchmark and programming languages implementations. The red line indicates the mean power draw of $189.8\pm 0.5$ W.
@@
Fundamentally, energy consumption is the product of power and time. Figure 9 decomposes these two factors. To test if the choice of programming language implementation has an effect on power draw beyond execution time, we fix other factors that may affect power draw. We normalize by execution time, and exclusively for this analysis, limit the program to a single core via Docker’s --cpuset-cpus argument. To ensure frequency scaling and throttling do not increase power draw variance in our measurements, we pin CPUs to their minimum frequency with turbo mode disabled. Figure 10 shows that there is no significant difference in power draw between programming language implementations once these outside factors are controlled for.
In theory, any language implementation with a runtime system running in parallel may affect energy consumption beyond variance in execution time, since increased core usage will increase power draw, as Section 4.4.1 details below. This hidden increased parallelism could occur with a parallel garbage collector or just-in-time compilation threads executing alongside the main program. In practice, compilation threads are primarily active only during startup, and both are rarely significant compared to the main program’s execution and other general overhead costs. For this reason, we draw this edge in gray in our model.
@@ 4.4. Contributors to Power Draw
(a) Average package (PKG) power draw by average number of active cores (§4.4.1). Log regression: $y=31\log_{2}x+246\left(R^{2}=0.72\right)$.
@@
(b) Average DRAM power draw by memory activity (§4.4.2), approximated using LLC misses. Linear regression: $y=1.68\cdot 10^{-8}x+12\left(R^{2}=0.93\right)$.
@@
Figure 12. Power draw is much more significantly affected by the number of active cores than memory activity (§4.4). Each point on those graphs represents a single benchmark run.
@@ 4.4.1. Number of Active Cores
Figure 12(a) shows the relationship between the number of active cores and power draw. Using multiple cores, energy usage grows up to around 2×2$\times$2 × the base power draw compared to using a single core. We fit a log curve to the data $\left(R^{2}=0.72\right)$, as the relationship does not appear linear, but results may vary across platforms. Denoting $\mbox{Power}\left(x\right)$ as the average power draw using $x$ cores, we find that, on our system, doubling the number of cores used increases average power draw by only roughly 31W:
$\mbox{Power}\left(x\right)$ $=31\log_{2}x+246$
$\Rightarrow\mbox{Power}\left(2x\right)$ $=31\log_{2}\left(2x\right)+246$
$\Rightarrow\mbox{Power}\left(2x\right)$ $=31\log_{2}2+31\log_{2}x+246$
$\Rightarrow\mbox{Power}\left(2x\right)$ $=\mbox{Power}\left(x\right)+31$
@@
$\frac{\mbox{Power}\left(2x\right)}{\mbox{Power}\left(x\right)}=\frac{31\log_{2}\left(2x\right)+246}{31\log_{2}x+246}\leq 113\%\quad\left(1\leq x\leq 128\right)$
@@ 4.4.2. Memory Activity
Previous research has shown that roughly 40% of dynamic random access memory (DRAM) energy consumption is constant and required to refresh memory cells to keep the system running properly, while the remaining 60% is related to read and write activity (Seol et al., 2016). Section 3.2.2 introduces last level cache (LLC) misses as a better proxy for memory activity: any read or write request that cannot be fulfilled by the LLC will have to go to DRAM. Figure 12(b) shows a strong linear correlation between LLC misses and increased power draw $\left(R^{2}=0.93\right)$.
@@ 4.5.1. Concurrency
Further, since TypeScript is a strict superset of JavaScript and we test them using the same runtime system, there should be no differences in execution time or energy consumption. With some minor edits in four benchmarks, all JavaScript benchmark implementations pass the TypeScript compiler without any errors, and yield equal performance.
@@ 4.5.2. Third-Party Libraries
Using third-party libraries also allows for easily crossing language boundaries, for instance to access lower-level system utilities or to obtain greater performance. In Python, using NumPy can yield up to 60,000×60,000$\times$60 , 000 × faster code when multiplying matrices (Leiserson et al., 2020; Berger et al., 2023). On a small micro-benchmark multiplying $000\times 10,$ 000 × 10 , 000 matrics, we find that NumPy is 84×84$\times$84 × faster than a naive three-loop C++ implementation, with $\times4$ lower energy consumption. Of course, this is because NumPy’s underlying matrix-multiply implementation is written in a low-level language and leverages many optimizations such as parallelism, vector instructions, blocking, or hardware-specific instructions. To some degree, using NumPy and other third-party libraries has become the “correct” way of writing Python code: leveraging low-level languages for compute-intensive tasks, while keeping Python’s simplicity. Even C++ may cross over to code originally written using C, Fortran, or direct Assembly for performance when using BLAS.
@@ 5. Related Work
Previous sections, notably Section 2, discuss the main line of research relating programming languages and energy consumption (Couto et al., 2017; Pereira et al., 2017, 2021). These studies discuss and rank programming languages based on execution time, energy consumption, and memory usage. Contrary to Pereira et al.’s results, we show in this paper that the interplay between these properties is complex and depends on many factors not taken into account in their work. Gordillo et al. (Gordillo et al., 2024) and Koedijk et al. (Koedijk and Oprescu, 2022) build on this line of research using direct hardware power measurements but suffer from similar limitations.
The processor’s operating voltage and frequency also have an impact on energy consumption. This impact has been studied in the context of trading off speed for reduced energy consumption, where compiling for speed yields better energy efficiency in the general case (Yuki and Rajopadhye, 2013). For some specific applications, such as sparse matrix computations, compile-time techniques may decrease energy usage without impact to execution time by leveraging load imbalance (Chen et al., 2005). Other work has shown that using characteristics of a program during execution to reduce the processor’s voltage or frequency can yield higher energy efficiency for a comparably smaller performance degradation (Hsu and Feng, 2005). For example, performance counter information can be used to scale the processor’s frequency (Weissel and Bellosa, 2002). These approaches are orthogonal to the choice of programming language and could be applied in many programming environments.
Graphics processing units (GPUs) are becoming increasingly prevalent in modern computing systems, notably in artificial intelligence workloads. Several studies have examined the energy efficiency of GPUs from a software perspective, notably accounting for the energy expenditure of tensors (Babakol and Liu, 2024) and comparing the energy consumption of different machine learning frameworks (Georgiou et al., 2022). Extending our work to incorporate a model for energy consumption of GPUs is a potential avenue for future work.
Previous research also addresses the energy consumption of different data structures in certain languages, notably Java collections (Hasan et al., 2016; Pereira et al., 2016; Oliveira et al., 2021), or Python data manipulation libraries (Nahrstedt et al., 2024). Data structures can be interesting candidates to study with an energy-focused lens as they can have a substantial impact on locality and memory activity, which are factors in energy consumption (as Section 4.4.2 describes). In many cases, however, programmers only have a limited choice of data structures that would be a good fit for their use case or algorithm.
Finally, past work has also investigated energy efficiency of concurrent programs, focusing on Haskell’s data sharing primitives (Lima et al., 2016) or Java’s thread management constructs (Pinto et al., 2014), or analyzing the power to performance tradeoffs in lock-based synchronization (Falsafi et al., 2016). As Section 4.4.1 argues, these studies confirm that the power usage patterns of the processor and the machine as a whole are more complex with parallel execution.
@@ 6. Conclusion
Using this causal model, we investigate and explain anomalies in previous research, finding that many factors such as parallelism level, benchmark implementation specifics, or language implementation properties must be taken into account for a fair comparison. Our results suggest that the choice of programming language has no significant impact on energy consumption beyond execution time. Programmers aiming to reduce energy consumption can do so even in “inefficient” programming languages like Python by using faster language implementations, employing faster and more parallel algorithms, and using fast third-party (native) libraries.